OpenStudy (anonymous):

Help! 3. Use the following functions. f(x) = 2x - 3 g(x) = 2 h(x) = 3x What is f(x)*g(x)

OpenStudy (anonymous):

Multiply (2x - 3)(2)

OpenStudy (michele_laino):

Please note that you have to compute this product: \[2\left( {x - 3} \right) \cdot 2 = ...?\]

OpenStudy (anonymous):

@Michele_Laino You multiplied by 2 twice.

OpenStudy (anonymous):

How would I multiply that?

OpenStudy (anonymous):

@iGreen. How would I multiply that?

OpenStudy (michele_laino):

Sorry, you are right @iGreen.

OpenStudy (bibby):

multiply term by term \(a(b-c)=(a*b)-(a*c)\)

OpenStudy (anonymous):

Would it be 4x-6?

OpenStudy (anonymous):

Yes, correct

OpenStudy (anonymous):

Okay. Thank you!

OpenStudy (anonymous):

Np